The probability that aa new automobile has a backup camera is 0.6. The probability



that a new automobile has a GPS system is 0.4. The probability that a new



automobile has both a backup camera and GPS is 0.2. If a new automobile is



selected at random, what is the probability that it has neither a backup camera nor a



GPS?



1
Expert's answer
2022-02-09T10:22:39-0500

Let "A" be the event that the new automobile has a backup camera. Its probability is "p(A)=0.6"

Let "B" be the event that the new automobile has a GPS. Its probability is "p(B)=0.4".

The probability that the new automobile has both a backup camera and GPS is "p(A\cap B)=0.2"

We first need to determine "p(A\cup B)=p(A)+p(B)-p(A\cap B)=0.6+0.4-0.2=0.8"

The probability that the new automobile has neither a backup camera nor a GPS is "1-0.8=0.2"

The required probability is 0.2